OpenStudy (anonymous):

NEED HELP WITH MATH PROBLEM! <3 Scott invested a total of $9500 at two separate banks. One bank pays simple interest of 12% per year while the other pays a simple interest at a rate of 7% per year. If Scott earned $960.00 in interest during a single year, how much did he have on deposit in each bank?

hero (hero):

x + y = 9500 .12x + .07y = 960

hero (hero):

x = amount in bank one y = amount in bank two

OpenStudy (anonymous):

okay thank you @Hero

hero (hero):

What did you get for x and y?

OpenStudy (anonymous):

im still working it out

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i already know my answer is wrong

OpenStudy (anonymous):

@Hero

hero (hero):

Multiply both sides of the second equation by 100 to get 12x + 7y = 96000 Multiply both sides of the first equation by 7 to get 7x + 7y = 7(9500) So now we have the following system: 12x + 7y = 96000 7x + 7y = 66500

hero (hero):

We can subtract the equations to eliminate y

hero (hero):

5x = 29500 x = 29500/5 x = 5900

hero (hero):

Try to find y